# A General Class of Model-Free Dense Precision Matrix Estimators

### v1

@@ 5.2 Methods
We conduct a comprehensive empirical analysis by fixing in-sample $n_{I}=240$ and considering both low-dimensional and high-dimensional cases across various portfolio sizes. Specifically, we selected 10, and 200 stocks from the pool to estimate the models with the same stocks for all windows in the low-dimensional experiments. For the high-dimensional experiment, we choose 357 (including all available stocks). We also analyzed portfolios with 25, 50, and 300 stocks, however, since these results are similar, they are not included here. We have 100 for low dimensional experiment and 325 stock high dimensional experiment in Online Appendix.
@@ 5.3 Results
We report the results for Maximum Sharpe Ratio portfolio, with and without transaction costs. We have also done analysis for Global Minimum Variance portfolio, and the results are very similar. Results for the low-dimensional cases are presented in Tables 1 - 2, while high-dimensional case is shown in Table 3. In additionally, we use the same data but with different out-of-sample from January 2005 to December 2023 ($n_{I}=120$, $n-n_{I}=228$). This subsample yields similar results and is available upon request. In MSR portfolio, our PCR-based method is the best model across all 2 small-to-medium (low-dimensional) portfolios scenarios for Sharpe Ratio with transaction cost. Our RRE achieves the highest Sharpe Ratio for the largest size of portfolio with $p=357$. Let us give an example by examining three portfolios of varying sizes for MSR: a small low-dimensional portfolio, a moderate-sized portfolio, and a large-sized portfolio. In Table 1, for a portfolio of 10 stocks ($p=10,n_{I}=240$), PCR-1F is the best Sharpe Ratio with 0.2260. In Table 2 of 200 stocks ($p=200,n_{I}=240$), the highest Sharpe Ratio is with our PCR-7F at 0.2682 and the second best is our PCR-3F at 0.2263. Lastly, in Table 3, for the largest portfolio with all 357 available stocks, the highest Sharpe ratio is RRE at 0.2203.
@@
Where $SR_{winner}$ is the method that has the highest Sharpe Ratio in each table, testing against all other methods and significance is set at threshold of $p=0.05$. A """""*"" ∗ " indicates statistical significance from the winner.
@@
Table 1: Monthly Portfolio Performance of 10 Stocks, $n_{I}=240$, $n-n_{I}=108$
@@
without transaction cost
RNW-GIC-SF 0.0101 0.0526 0.1913 0.0953
RNW-GIC-3F 0.0101 0.0519 0.1945 0.0889
RNW-GIC-5F 0.0102 0.0510 0.2002 0.0876
NW-GIC 0.0106 0.0505 0.2093 0.0888
POET 0.0104 0.0507 0.2051 0.0700
Ledoit-Wolf 0.0100 0.0522 0.1914 0.0927
Inter-RRE 0.0104 0.0530 0.1957 0.0929
PCR-1F 0.0109 0.0490 0.2221 0.0518
PCR-3F 0.0105 0.0492 0.2142 0.0576
PCR-5F 0.0106 0.0501 0.2125 0.0948
PCR-7F 0.0105 0.0507 0.2075 0.0795
PCR-Adaptive 0.0108 0.0489 0.2211 0.0514
@@
Table 2: Monthly Portfolio Performance of 200 Stocks, $n_{I}=240$, $n-n_{I}=108$
@@
without transaction cost
RNW-GIC-SF 0.0053 0.0442 0.1205 0.2761
RNW-GIC-3F 0.0067 0.0447 0.1496 0.3171
RNW-GIC-5F 0.0072 0.0484 0.1483 0.3996
NW-GIC 0.0093 0.0429 0.2169 0.1517
POET 0.0057 0.0430 0.1336 0.1989
Ledoit-Wolf 0.0048 0.0521 0.0928 0.7320
Inter-RRE -0.0018 0.1168 -0.0158 6.1638
PCR-1F 0.0075 0.0376 0.2002 0.1203
PCR-3F 0.0082 0.0382 0.2151 0.2860
PCR-5F 0.0081 0.0557 0.1462 1.0298
PCR-7F 0.0128 0.0433 0.2964 1.0439
PCR-Adaptive 0.0081 0.0435 0.1865 0.4231
@@
Table 3: Monthly Portfolio Performance of 357 Stocks, $n_{I}=240$, $n-n_{I}=108$
@@
without transaction cost
RNW-GIC-SF 0.0046 0.0432 0.1068 0.2857
RNW-GIC-3F 0.0057 0.0440 0.1297 0.3580
RNW-GIC-5F 0.0056 0.0461 0.1209 0.4507
NW-GIC 0.0093 0.0422 0.2202 0.1333
POET 0.0058 0.0424 0.1364 0.2009
Ledoit-Wolf 0.0015 0.0534 0.0280 0.9073
Inter-RRE 0.0106 0.0453 0.2338 0.2725
PCR-1F 0.0073 0.0364 0.2004 0.1349
PCR-3F 0.0054 0.0366 0.1470 0.2798
PCR-5F 0.0056 0.0528 0.1056 1.6010
PCR-7F 0.0095 0.0506 0.1870 1.8931
PCR-Adaptive 0.0085 0.0420 0.2020 0.2608
@@ 6 Conclusion
In this paper, we have introduced a general precision matrix estimator. Our estimator is model-free and does not rely on any sparsity assumptions. Under certain conditions, our estimator exhibits double descent in the estimation error. A distinguishing feature of our analysis is that it is done for finite samples, and leads to non-asymptotic bounds for the error of precision matrix estimator. Our numerical results support our theoretical findings. We have shown that the proposed estimator compares favorably with popular existing machine learning estimator alternatives. Moreover, the Sharpe ratio produced by one of our estimators on the S&P 500 data exhibits the double ascent, which to the best of our knowledge has never obtained in other empirical studies.

### v2

@@ 5.2 Methods
We conduct a comprehensive empirical analysis by fixing in-sample $n_{I}=240$ and considering both low-dimensional and high-dimensional cases across various portfolio sizes. Specifically, we selected 10, and 200 stocks from the pool to estimate the models with the same stocks for all windows in the low-dimensional experiments. For the high-dimensional experiment, we choose 357 (including all available stocks). We also analyzed portfolios with 25, 50, and 300 stocks, however, since these results are similar, they are not included here.
@@ 5.3 Results
We report the results for Maximum Sharpe Ratio portfolio, with transaction costs in Tables 1-3, and $NA$ represents ”not available” for turnover in our tables with transaction costs. $SD$ represents Standard Deviation of the portfolio, and $Sharpe$ represents Sharpe Ratio of the portfolio respectively in Tables.
Tables without transaction costs are in Appendix and also have turnover results. We have also done analysis for Global Minimum Variance portfolio, and the results are very similar. Results for the low-dimensional cases are presented in Tables C.2 - C.3, while high-dimensional case is shown in Table C.4. In additionally, we use the same data but with different out-of-sample from January 2005 to December 2023 ($_{I}=120, nn$$n_{I}=228$). This subsample yields similar results and is available upon request. In MSR portfolio, our PCR-based method is the best model across all 2 small-to-medium (low-dimensional) portfolios scenarios for Sharpe Ratio with transaction cost. Our RRE achieves the highest Sharpe Ratio for the largest size of portfolio with $p=357$. Let us give an example by examining three portfolios of varying sizes for MSR: a small low-dimensional portfolio, a moderate-sized portfolio, and a large-sized portfolio. In Table C.2, for a portfolio of 10 stocks ($p=10,n_{I}=240$), PCR-1F is the best Sharpe Ratio with 0.2260. In Table C.3 of 200 stocks ($p=200,n_{I}=240$), the highest Sharpe Ratio is with our PCR-7F at 0.2682 and the second best is our PCR-3F at 0.2263. Lastly, in Table C.4, for the largest portfolio with all 357 available stocks, the highest Sharpe ratio is RRE at 0.2203.
@@
Where $SR_{winner}$ is the method that has the highest Sharpe Ratio in each table, testing against all other methods and significance is set at threshold of $p=0.05$. A "*" indicates statistical significance from the winner.
@@
Table 1: Monthly Portfolio Performance of 10 Stocks, $_{I}=240, nn$$n_{I}=108$
@@
Table 2: Monthly Portfolio Performance of 200 Stocks, $_{I}=240, nn$$n_{I}=108$
@@
Table 3: Monthly Portfolio Performance of 357 Stocks, $_{I}=240, nn$$n_{I}=108$
@@ 6 Conclusion
In this paper, we have introduced a general precision matrix estimator. Our estimator is model-free and does not rely on any sparsity assumptions. Under certain conditions, our estimator exhibits double descent in the estimation error. Moreover, the Sharpe ratio produced by one of our estimators on the S&P 500 data exhibits the double ascent, which to the best of our knowledge has never obtained in other empirical studies.